Circular No. 6982

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998di IN ANONYMOUS GALAXY M. Schwartz, Cottage Grove, OR, reports his discovery of an apparent

supernova (mag 16.8) on an unfiltered CCD frame taken with the 0.35m Tenagra I automated supernova patrol telescope on Aug. 2.48 UT. SN 1998di is located at α = 19^{h}24^{m}38^{s}.19, δ = +59^{o}41′46′′.7 (equinox 2000.0), which is ∼ 0′′.6 east and 3′′.8 north of the nucleus of an anonymous galaxy. The new object was confirmed on a frame taken on Aug. 3.26. No star is present at this position on a frame taken on July 22.46 using the same equipment (limiting mag ∼ 19). Schwartz adds that the object does not appear on either the first or the second Palomar Sky Survey scans, and that it is absent from the USNO A1.0 catalogue.

CATACLYSMIC VARIABLE IN PEGASUS The following unfiltered CCD magnitudes of this variable star (cf.

IAUC 6746, 6772), reported by Q.-Y. Qiao, J.-Y. Wei, Y.-l. Qiu, and J.-Y. Hu, Beijing Astronomical Observatory, were obtained with the 0.60-m telescope at Xinglong Station: July 19.75 UT, > 18.0; 28.73, 15.6; 29.62, 16.0. On July 29.77 they measured R = 15.8 (comparison-star magnitudes from USNO A1.0 catalogue).

SUPERNOVAE 1998an, 1998bn, 1998bp, 1998bu, 1998cs CCD magnitudes (unfiltered unless otherwise noted; B = L. Boschini,

Monte di Nese, Bergamo, Italy; H = D. Hanžl, Brno, Czech Republic; M = G. Masi, Ceccano, Italy; T = S. Moretti and S. Tomaselli, Forlí, Italy):

SN 1998an in UGC 3683: Apr. 20.835 UT, V = 16.17 (H); 25.841, V = 16.37 ± 0.07 (H).

SN 1998bn in NGC 4462: May 2.883 UT, V = 13.39 ± 0.02, B − V = +0.54 ± 0.06 (H); 5.864, V = 13.42 ± 0.04, B − V = +0.67 ± 0.04 (H); 8.841, V = 13.53 ± 0.01, B − V = +0.67 ± 0.05 (H); 14.85, 15.0 (T).

SN 1998bp in NGC 6495: Apr. 30.980 UT, R = 14.8 (H); May 8.948, V = 14.74 (H); 9.11, 14.6 (B); 10.087, 14.9 (M); 11.94, 15.1 (T).

SN 1998bu in NGC 3368 (cf. IAUC 6937): May 16.883 UT, 12.1 (M); 22.827, 12.0 (M); June 7.888, 12.7 (M); 19.855, V = 13.52 ± 0.04, B − V = +1.4 ± 0.5 (H); 20.877, V = 13.54 ± 0.06 (H); 21.862, V = 13.63 ± 0.01 (H); 25.874, V = 13.75 ± 0.05 (H); 27.859, V = 13.90 ± 0.08 (H).

SN 1998cs in UGC 10432: July 2.95 UT, 17.1 (T).

1998 August 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green